OpenStudy (jenniferjuice):

simplify 2/2+5i 2/29 4/29 4+10i/29 4_10i/29

OpenStudy (anonymous):

\(\Large \frac{2}{2+5i}=\frac{2(2-5i)}{(2+5i)(2-5i)}=\frac{4-10i}{4-25i^2}\) \(\Large i^2=-1 \Rightarrow \frac{4-10i}{4-25i^2}=\frac{4-10i}{4+25}=\frac{4-10i}{29}\)